vapour pressure of 4% solution of a nonvolatile solute in water at 100 centigrade is 745mm .what is the molecular weight of solute

Grade:

3 Answers

Sunil Kumar FP
9 years ago
760-745/760=4*18/M*96
M=38
Sunil Kumar FP
9 years ago
760-745/760=4*18/M*96
M=38
Sunil Kumar FP
9 years ago
thanks and regards
sunil kr
askIITiab faculty